# Slice-Level Scheduling for High Throughput and Load Balanced LLM Serving

### v1

@@ Abstract.
Large language models (LLMs) iteratively generate text token by token, with memory usage increasing with the length of generated token sequences. The unpredictability of generation lengths makes it difficult to estimate the time and memory needed to process requests, posing a challenge for effective request scheduling. Conventional sequence-level scheduling (SLS) serves requests in a first-come first-served (FCFS) manner with static batching where requests with short generation lengths are delayed until those with long ones have finished generation, which hurts computational efficiency. Besides, to avoid out-of-memory (OOM) errors, SLS batches requests with a small batch size, which limits throughput. Recently proposed iteration-level scheduling (ILS) enhances computational efficiency with continuous batching to return completed requests timely and dynamically add new requests for processing. However, many ILS schedulers limit the number of parallel-processing requests to avoid OOM errors while achieving a fast inference speed, which compromises throughput. Moreover, existing SLS and ILS schedulers fail to balance the workload across multiple deployed LLM instances. To tackle these challenges, we propose slice-level scheduling (SCLS). By splitting the predefined maximal generation length limit into slices and serving batches slice by slice, it provides a precise range of serving time and memory usage for batched requests, laying the foundation for effective scheduling. Experiments confirm that compared with SLS and ILS schedulers, SCLS can improve throughput by up to 315.8% and greatly mitigate load imbalance with proposed batching and offloading algorithms.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
@@
With increasing the parameter size, generative large language models (LLMs) have demonstrated remarkable capabilities across a wide range of SQL-related tasks, such as natural language to SQL, SQL interpretation, and SQL optimization. These advancements greatly benefit LLM-based conversational code-assistant agents like CodeFuse (codefuse, ) developed by Ant Group. Besides SQL, CodeFuse is also widely acclaimed for its ability to generate, optimize, and interpret various programming languages including Java, Python, and C++, boasting a peak concurrent usage with thousands of online users. Therefore, an efficient serving system with high throughput and scalability is essential to handle such a heavy workload.
LLMs generate tokens iteratively, one per iteration until the terminator end-of-sequence (EOS) token is generated. The number of iterations required to generate the EOS token is referred to as the request generation length. To prevent the generation process from being too long, a maximal generation length limit is always set. When this limit is reached, even if the EOS token has not been generated, the generated results will be returned to users. Moreover, due to the key-value cache that is always leveraged to accelerate LLM inference, the GPU memory consumption increases with the increased number of generated tokens. Therefore, the generation length of a request determines the computational overhead and the memory it consumes in the generation process.
@@
To achieve high throughput for LLM serving, it is significant to fully utilize the parallel computing capability of GPUs with batch serving. Static batching and continuous batching are two batch-serving techniques for LLMs. Static batching pads the raw input of batched requests to the same length before serving and iteratively generates tokens until all requests generate the EOS token or the predefined iteration number limit is reached. Continuous batching leverages customized CUDA kernels to remove request padding and enable dynamic exits and joins of requests during batch serving. However, since the request generation length is unpredictable, it is difficult to estimate the computational overhead and memory consumption of requests. Hence, it is challenging to schedule requests to take full advantage of these batch-serving techniques.
@@
Figure 1. Illustration of various scheduling techniques, where $r_{i}$ represents the $i$th arrived request.
Existing deep learning serving systems (triton, ; tensorflowserving, ) adopt the sequence-level scheduling (SLS) to batch requests in a first-come first-served (FCFS) manner with a fixed batch size and serve them with static batching as shown in Fig. 1(a), which leads to severe computational inefficiencies. Firstly, without knowing request generation lengths, the SLS scheduler directly sets the iteration number limit for static batching to the predefined maximal generation length and defaults to a small batch size to prevent out-of-memory (OOM) errors, thus failing to fully exploit the GPU’s parallel computing capability and compromising throughput. Secondly, in a batch, requests with short inputs are padded and requests with short outputs have to wait for those with long outputs to finish before they are returned together. Besides, during the waiting, completed requests still participate in the computation and generate invalid tokens.
To eliminate the request padding and waiting caused by SLS and improve computational efficiency, iteration-level scheduling (ILS) is introduced in recent studies (yu2022orca, ; kwon2023efficient, ; deepspeed-fastgen, ). Current ILS techniques leverage continuous batching to timely return completed requests and add newly arrived requests for processing in an FCFS manner at each iteration as shown in Fig. 1(b). However, without knowing request generation lengths, many ILS techniques (yu2022orca, ; deepspeed-fastgen, ) adopt conservative memory management strategies that limit the number of parallel-processing requests to avoid OOM errors and speed up inference, thus under-utilizing the GPU and hurting throughput.
Moreover, existing SLS and ILS schedulers (triton, ; tensorflowserving, ; deepspeed-fastgen, ) leverage the round-robin policy to offload requests to multiple LLM instances, which can cause unbalanced distribution of computational overhead and memory consumption across LLM instances due to the variance of request generation lengths.
@@
In this paper, we propose slice-level scheduling (SCLS) to address the aforementioned challenges from a pure scheduling perspective. The core idea is to split the predefined maximal generation length limit into fixed-length slices and serve batches slice by slice. Requests that complete their generation are returned, while incomplete requests are rescheduled with newly arrived requests until finished. SCLS can integrate with both static and continuous batching. We focus on the integration with static batching in this paper and discuss the integration with continuous batching in Section 7.
@@
By setting the iteration number limit of static batching to the slice length, requests with short generation lengths can be returned in a few slices without waiting for long. Besides, compared with the predefined maximal generation length, the small slice length results in a narrow range of generation length at each schedule, thus providing a precise range of memory consumption and serving time for batches. Therefore, with the precise range of memory consumption, requests can be adaptively batched together with as large a batch size as possible to improve the throughput without exceeding GPU memory as shown in Fig. 1(c). Moreover, with the precise range of serving time, batches can be scheduled to LLM instances using the max-min policy (radunovic2007unified, ) to achieve load-balancing. All in all, generation slicing lays the foundation for effective scheduling.
@@
We implement a prototype system of SCLS and conduct extensive experiments with 8 LLaMA2-13B instances deployed on 8 NVIDIA A100 80GB GPUs. The experimental results show that SCLS can achieve a high request throughput and a more balanced workload across LLM instances. Our main contribution is listed as follows.
@@
  • We demonstrate the inefficiency and load imbalance of existing SLS and ILS schedulers through experiments. Besides, we find that requests with long outputs are rare by analyzing the ShareGPT dataset and request traces collected from CodeFuse logs. These motivate the birth of SCLS.
@@
  • We conduct extensive experiments to confirm the superiority of SCLS. Equipped with various inference engines, SCLS can improve request throughput by up to 315.8% and reduce response time by up to 91.1% compared with existing SLS and ILS schedulers under the CodeFuse request trace.
2. Preliminaries
2.1. Autoregressive Language Model Generation
Language models take a token sequence as input to predict the probability distribution over the vocabulary for the next token. By sampling from the distribution, a new token is generated. Hence, a sequence of tokens can be generated by iteratively adding the newly generated token at the end of the input sequence and re-feeding the sequence to the language model. Such a generation pattern is called autoregressive generation.
The autoregressive generation stops when the EOS token is generated or the predefined maximal generation length limit is reached. In this paper, the process of generating one token is termed an iteration. The length (i.e., number of tokens) of the raw request text is termed as the request input length. Besides, we term the number of iterations until the EOS token is output as the request generation length.
Figure 2. LLM inference procedure. The blue and green rounded rectangles depict the computation flow of the transformer block and masked-self attention, respectively.
2.2. Large Language Model Inference Procedure
Current popular LLMs such as GPT, LLaMA, LLaMA2, Qwen, and ChatGLM (brown2020language, ; touvron2023llama, ; touvron2023llama2, ; du2022glm, ; bai2023qwen, ; zeng2022glm, ) all adopt the decoder-only transformer architecture whose computation flow is depicted in Fig. 2.
After feeding the text ”I want to” to the LLM, the input token sequence is successively processed by a linear layer called embedding, a stack of transformer blocks, a linear layer called lm-head, and finally a possibility distribution is produced, from which ”study” is sampled as the next token.
The most significant part of LLMs is the transformer block, of which the masked self-attention module is the core component that elevates it (vaswani2017attention, ). As shown in Fig. 2, for each token in the input sequence, the self-attention module first derives three tensors, which are query, key, and value. After that, the query matrix and key matrix are multiplied to yield the attention-score matrix, which represents the correlation between tokens. Since LLMs are trained to predict the next token, each input token should not see information after its location. Therefore, a mask operation is performed on the attention-score matrix, where a very large negative number is added to the attention scores that represent the correlation between tokens and their successive tokens. Therefore, these attention scores become negligible in the relevant matrix obtained after the soft-max operation. Next, by multiplying the relevant matrix with the value matrix, each token can extract features from its preceding tokens to update itself.
2.3. Key-Value Cache and Two-Phase Inference
In a stateless implementation of LLM autoregressive generation, each transformer block computes key and value tensors for all input tokens in each iteration, resulting in significant computational overhead. This overhead escalates in autoregressive generation because newly generated tokens are continuously appended to the input sequence, which incrementally increases the per-iteration computational overhead.
Figure 3. Key-value cache usage in two phases. The raw request has 3 tokens and the blue and green grids represent the newly derived and reused key and value tensors, respectively.
@@ 2.3. Key-Value Cache and Two-Phase Inference
To alleviate such a severe overhead, existing inference engines such as huggingface-transformers (huggingface-transformers, ) and deepspeed-inference (deepspeed-inference, ) cache the computed key and value tensors in GPU memory. As shown in Fig. 3, in the first iteration, the origin input sequence is fed to the LLM and the corresponding key and value tensors are computed and cached. In subsequent iterations, only the latest generated token is fed to the LLM for computation and cache of key and value tensors, which significantly cuts down the computational overhead. Accordingly, the inference process can be split into two phases. The first iteration is the prefill phase, which involves intensive key and value computation for all raw input tokens of the request, hence computationally heavy. Subsequent iterations are the decoding phase, which only computes key and value tensors for the input token, hence computationally light.
Figure 4. Static batching example. The blue and green grids represent raw input tokens of the requests and valid tokens generated by the LLM, respectively.
2.4. Serving Requests with Static Batching
Fig. 4 illustrates the serving process of static batching. Firstly, all the batched requests are padded to the same length as the longest one using the pad token. Next, the batch is fed to the LLM to generate tokens autoregressively. To prevent the pad token from being attended by other tokens, the corresponding attention score is also masked during the inference process. Until all the requests generate the EOS token or the preset iteration number limit is reached, the serving process completes. Since the generation lengths of requests in a batch are diverse, the requests that generate the EOS token early continue generating invalid tokens until the serving process completes, but the tokens generated after the EOS token will be ignored in the response.
In this paper, we refer to the longest raw input length of requests in a batch as the batch input length, which is equal to the token sequence length of each request after padding. Moreover, we take the minimum of the preset iteration number limit and the longest generation length of requests in a batch as the batch generation length, which is equal to the total number of iterations in the batch serving procedure.
@@ 2.4. Serving Requests with Static Batching
(c) Pad token number.
(d) Invalid token number.
@@
Figure 5. Inefficiency and load imbalance of SLS and ILS schedulers. Sub-figures (a) and (e) show that SCLS can significantly outperform SLS and ILS schedulers in terms of throughput and load balancing. Sub-figures (b)~(d) present that SCLS’s superior performance comes from the increase of batch size and decrease of the number of pad tokens and invalid tokens.
3. Motivation
3.1. Inefficient Serving of SLS and ILS
We conduct an experiment to confirm the inefficient serving of SLS and ILS schedulers. We implement the SLS and SCLS schedulers using the deepspeed-inference-v0.13.3 as the inference engine and leverage the deepspeed-fastgen-v0.2.0 as the ILS scheduler. In the experiment, 8 LLaMA2-13B instances are deployed on 8 NVIDIA A100 80GB GPUs, and requests with various input and generation lengths arrive at a rate of 20 per second. Other settings are the same as in Section 5. The experimental results are shown in Fig. 5.
From Fig. 5(a), we can see that the throughput of SLS is much smaller than that of SCLS. This is because SLS serves requests with a fixed small batch size, which wastes the powerful parallel computing capability of GPUs. Moreover, SLS batches requests in an FCFS manner, which leads to a lot of pad tokens and incurs additional computational overhead. Furthermore, setting the iteration number limit of static batching directly to the maximal generation length limit causes many requests can not be returned in time and a large number of invalid tokens are generated.
Although the continuous batching can eliminate pad tokens and invalid tokens in the serving process, the throughput of ILS is also much lower than that of SCLS, because the ILS scheduler adopts a conservative memory management mechanism that limits the number of parallel-processing requests. Therefore, the ILS scheduler can not fully utilize the parallel computing capability of GPUs, resulting in a low throughput.
3.2. Unbalanced Workload of SLS and ILS
We record the time when each LLM instance completes serving at the end of the experiment and depict the standard deviation (STD) of the completion time (CT) of instances in Fig. 5(e).
Since both SLS and ILS schedulers use the round-robin policy to offload requests, requests with long generation lengths and requests with short generation lengths may be assigned to different instances, resulting in unbalanced memory consumption and computational overhead across instances. For the SLS scheduler, instances that have more requests with long generation lengths take more time to serve requests. For the ILS scheduler, instances that have more requests with long generation lengths consume more memory and cannot often add new requests for processing due to insufficient memory, causing a long request queuing time. Such a load imbalance accumulates to result in a large variance in the completion time of LLM instances. Hence, the standard deviation of instance completion time of SLS and ILS is much larger than that of SCLS.
@@ 3.2. Unbalanced Workload of SLS and ILS
Figure 6. Probability density function (PDF) and cumulative distribution function (CDF) of the request generation length in actual scenarios.
3.3. Actual Generation Length Distribution
To understand the generation length distribution in actual scenarios, we analyze the request data of CodeFuse and ShareGPT. CodeFuse (codefuse, ) is a chatbot developed by Ant Group that is optimized for generating, interpreting, and optimizing program code. We collected user request trace from CodeFuse’s logs for October and November 2023, and depict the distribution of request generation lengths in Fig. 6(a). ShareGPT (sharegpt, ) is a publicly available website that shares conversations from users chatting with ChatGPT (chatgpt, ). We analyze close to 400,000 pieces of data collected from ShareGPT and present the generation length distribution in Fig. 6(b).
From Fig. 6, we can find that the vast majority of requests have a small generation length of less than 512 in both the CodeFuse and ShareGPT data. Hence, for static batching, if we split the maximal generation length limit (e.g., 2048) into slices with a small slice length (e.g., 128) and serve batched requests slice by slice, most requests can be returned in time within a small number of slices. In addition, since the iteration number limit is set to the small slice length, we can get a precise range of memory consumption and batch together as many uncompleted requests and newly arrived requests as possible while ensuring that OOM errors will not occur, thus increasing throughput and achieving efficient batch serving.
Moreover, we can also get a precise range of batch serving time as well. Therefore, load balancing can be achieved by offloading batches with the max-min policy where the batch with the longest serving time is offloaded to the least loaded instance.
Considering the above, we propose SCLS, which splits the predefined maximal generation length limit into fixed-length slices, limits the number of iterations in each batch serving to the slice length and serves requests slice by slice. Although requests with long generation lengths suffer from the extra overhead of recomputing the prefill phase at every reschedule, the overhead is acceptable because these long-generated requests are rare as shown in Fig. 6.

### v2

# Slice-Level Scheduling for High-Throughput and Load-Balanced LLM Serving

@@ Abstract
Large language models (LLMs) iteratively generate text token by token, with memory usage increasing with the length of generated token sequences. Since the request generation length is generally unpredictable, it is difficult to estimate the time and memory required to process requests, thus posing a challenge for effective request scheduling. Conventional sequence-level scheduling (SLS) serves requests in a first-come first-served (FCFS) manner with static batching where requests with short generation lengths are delayed until those with long ones have finished generation. Besides, to avoid out-of-memory (OOM) errors, SLS batches requests using a small batch size, which limits throughput. Recently proposed iteration-level scheduling (ILS) improves this with continuous batching, timely completing requests and dynamically adding new ones, but often limits the number of parallel-processing requests to OOM errors, thus compromising throughput. Moreover, both SLS and ILS fail to effectively balance workload across multiple LLM instances. To tackle these challenges, we propose slice-level scheduling (SCLS). By splitting the predefined maximal generation length limit into slices and serving batches slice by slice, it provides a precise range of serving time and memory usage for batched requests, laying the foundation for effective scheduling. Experiments confirm that compared with SLS and ILS schedulers, SCLS can improve throughput by up to 315.8% and greatly mitigate load imbalance with proposed batching and offloading algorithms.
Index Terms:
Large language model, machine learning system, transformer serving, adaptive batching, load balancing.
I Introduction
@@
Large language models (LLMs) are currently playing an important role in a wide range of applications, especially interactive web chatbots, such as CodeFuse [1] developed by Ant Group. In addition to the daily chat capability, CodeFuse is also widely acclaimed for its enhanced ability to generate, optimize, and interpret various programming languages, including SQL, Java, Python, and C++, boasting a peak concurrent usage with thousands of online users. Therefore, an efficient serving system with high throughput and scalability is essential to handle such a heavy workload.
@@
To achieve high throughput for LLM serving, it is significant to fully utilize the parallel computing capability of GPUs with batch serving. Static batching and continuous batching are two batch-serving techniques for LLM inference. Static batching pads the raw input of batched requests to the same length before serving and iteratively generates tokens until all requests generate the end-of-sequence (EOS) token or the predefined iteration number limit is reached. Continuous batching leverages customized CUDA kernels to remove request padding and enable dynamic exits and joins of requests during batch serving.
However, the time and GPU memory required for serving a request increases with the generation length (i.e., the number of generated tokens) because of the autoregressive generation pattern and key-value cache mechanism of LLM inference. Since the request generation length is generally unpredictable, it is difficult to estimate the serving time and memory consumption of requests in advance. Therefore, it is challenging to schedule requests to take full advantage of these batch-serving techniques for high throughput.
Existing deep learning serving systems [2, 3] adopt the sequence-level scheduling (SLS) to batch requests in a first-come first-served (FCFS) manner with a fixed batch size and serve them with static batching as shown in Fig. 1(a), which leads to severe computational inefficiencies. Firstly, for static batching, in a batch, requests with short inputs are padded, and requests with short outputs have to wait for those with long outputs to finish before they are returned together. During the waiting, completed requests still participate in the computation and generate invalid tokens. Secondly, without knowing request generation lengths, the SLS scheduler directly sets the iteration number limit for static batching to the predefined maximal generation length and defaults to a small batch size to prevent out-of-memory (OOM) errors, thus failing to fully exploit the GPU’s parallel computing capability and compromising throughput.
To eliminate the request padding and waiting caused by SLS and improve computational efficiency, iteration-level scheduling (ILS) is introduced in recent studies [4, 5, 6]. Current ILS techniques leverage continuous batching to timely return completed requests and add newly arrived requests for processing in an FCFS manner at each iteration, as shown in Fig. 1(b). However, without knowing request generation lengths, many ILS techniques [4, 6] adopt conservative memory management strategies that limit the number of parallel-processing requests to avoid OOM errors, thus under-utilizing the GPU and hurting throughput.
Moreover, existing SLS and ILS schedulers [2, 3, 6] leverage the round-robin policy to offload requests to multiple LLM instances, which causes unbalanced distribution of computational overhead and memory consumption across LLM instances due to the variance of request generation lengths.
@@
Figure 1: Illustration of various scheduling techniques, where $r_{i}$ represents the $i$th arrived request.
@@
In this paper, we propose slice-level scheduling (SCLS) to address the aforementioned challenges from a pure scheduling perspective. The core idea is to split the predefined maximal generation length limit into fixed-length slices and serve batches slice by slice. Completed requests are returned promptly, while incomplete requests are rescheduled with newly arrived requests until finished. Generation slicing can enhance both static and continuous batching. We focus on strengthening static batching in this paper. In Section VII, we discuss promising technique solutions for applying the idea of generation slicing to continuous batching, which we leave as our future work.
@@
Compared with the maximal generation length limit, the small slice length results in a narrow range of generation length at each schedule, thus providing a precise range of memory consumption and serving time for batches. With the precise range of memory consumption, requests can be adaptively batched together with as large a batch size as possible to improve the throughput without exceeding the GPU memory as shown in Fig. 1(c). With the precise range of serving time, batches can be scheduled to LLM instances using the max-min policy [7] to achieve load-balancing. All in all, generation slicing lays the foundation for effective scheduling. Besides, since requests with long outputs are rare in real-world applications such as ChatGPT [8] and CodeFuse, most requests can be returned in a few slices without waiting for long.
@@
We implement SCLS as the gateway to route requests and conduct extensive experiments with LLM instances deployed in the production environment of Ant Group using CodeFuse request traces. Experimental results show that SCLS can achieve a high request throughput and a more balanced workload across LLM instances. Our main contribution is listed as follows.
@@
  • We confirm the inefficiency and load imbalance of SLS and ILS schedulers through experiments. Besides, we find that requests with long outputs are rare by analyzing the ShareGPT dataset and request traces collected from CodeFuse’s logs. These motivate the birth of SCLS.
@@
  • We conduct extensive experiments to confirm the superiority of SCLS. Equipped with various inference engines, SCLS can improve request throughput by up to 315.8%, reduce response time by up to 91.1%, and greatly mitigate load imbalance compared with SLS and ILS schedulers.
II Preliminaries
LLMs generate text by predicting the next token for an input token sequence and adding the predicted token to the input iteratively until the EOS token is generated. Such a generation pattern is called auto-regressive generation. To prevent the generation process from being too long, a maximal generation length limit is always set. When this limit is reached, even if the EOS token has not been generated, the generated results are returned to users. The request input length is the number of tokens in the raw request text, while the request generation length is the number of generated tokens.
In static batching, requests are padded to the same length and processed together. Therefore, the batch input length is the length of the longest request in the batch. The batch generation length is the total number of iterations for the batch, depending on the longest request generation length and the iteration number limit.
@@ II Preliminaries
To accelerate LLM inference, the key and value tensors produced by transformer blocks of LLMs are cached for reuse. In the first iteration of the auto-regressive generation, the token sequence of raw request text is fed to the LLM, and the key and value tensors of these input tokens are computed and cached. In subsequent iterations, only the latest generated token is fed to the LLM for computation and cache of key and value tensors. Thus, the generation process can be split into two phases. The first iteration is the computationally heavy prefill phase. Subsequent iterations are the computationally light decoding phase.
@@
(c) #Pad token.
(d) #Invalid token.
@@
Figure 2: Inefficiency and load imbalance of SLS and ILS schedulers. Sub-figures (a) and (e) show that SCLS can significantly outperform SLS and ILS schedulers in terms of throughput and load balancing. Sub-figures (b)~(d) present that SCLS’s superior performance comes from the increase of batch size and decrease of the number of pad tokens and invalid tokens.
III Motivation
III-A Inefficient Serving of SLS and ILS
We conduct experiments to confirm the inefficient serving of SLS and ILS schedulers. We implement the SLS and SCLS schedulers using deepspeed-inference [9] as the inference engine and leverage deepspeed-fastgen [6] as the ILS scheduler. In the experiment, 8 LLaMA2-13B [10] instances are deployed on 8 NVIDIA A100 80GB GPUs, and requests with various input and generation lengths arrive at a rate of 20 per second. Other settings are the same as in Section V. The experimental results are shown in Fig. 2.
From Fig. 2(a), we can see that the throughput of SLS is much smaller than that of SCLS. This is because SLS serves requests with a fixed small batch size. Moreover, SLS simply batches requests in an FCFS manner, which leads to a lot of pad tokens and generated invalid tokens caused by the variance of the batched requests’ input lengths and generation lengths. Besides, although continuous batching can eliminate pad tokens and invalid tokens in the serving process, the throughput of ILS is also much lower than that of SCLS, because the ILS scheduler adopts a conservative memory management mechanism that limits the number of parallel-processing requests.
@@ III-A Inefficient Serving of SLS and ILS
Figure 3: Probability density function (PDF) and cumulative distribution function (CDF) of the request generation length in actual scenarios.
III-B Unbalanced Workload of SLS and ILS
We record the time when each LLM instance completes serving at the end of the experiment. Fig. 2(e) presents the standard deviation (STD) of the completion time (CT) of LLM instances. Since both SLS and ILS schedulers use the round-robin policy to offload requests, requests with long and short generation lengths may be assigned to different instances, resulting in unbalanced memory consumption and computational overhead across instances. For the SLS scheduler, instances that have more requests with long generation lengths take more time to serve requests. For the ILS scheduler, instances that have more requests with long generation lengths consume more memory and cannot promptly add new requests for processing due to insufficient memory, causing a long request queuing time. Such a load imbalance accumulates to result in a large variance in the CT of LLM instances. Hence, the CT STD for LLM instances under SLS and ILS schedulers is much larger than that of SCLS.